OpenStudy (anonymous):

What is the value of the expression 8–^2?

OpenStudy (mathmale):

Your meaning here is unclear. Please use either Equation Editor (below) or the Draw utility (below) in sharing your expression.

OpenStudy (anonymous):

I think you mean 8^(-2), in which case, 8^(-2) = 1/8^2 = 1/64